An investor spends a total of $15,000 in two investments, a mortgage and a bond. The mortgage pays 10% interest and the bond pays 6% interest. If the total interest from both investments for one year is $1,120 find the amount invested in the mortgage and the bond.
Thank You!!

Answer by checkley77(12844)

.10x+.06(15,000-x)=1,120
.10x+900-.06x=1,120
.04x=1,120-900
.04x=220
x=220/.04
x=5,500 amount invested @ 10%
15,000-5,500=9,500 invested @ 6%
Proof:
.10*5,500+.06*9,500=1,120
550+570=1,120
1,120=1,120